OpenStudy (anonymous):

Is there only one correct way to solve a multi-step equation containing grouping symbols and variables on both sides? Explain. Use example in your answer

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

there's multiple ways to solve a multi step equation, except I can't really explain it :P I just know this because I solve it way differently than the teacher teaches but my way is faster and easier :I

OpenStudy (anonymous):

My teacher asked me this but I don't know what he is talking about?

OpenStudy (skullpatrol):

What do you think he is talking about?

OpenStudy (anonymous):

He is talking about a equation that looks like 3x + 4 =10 - 5x with the variable on both sides of the equal sign the answer to the question is x=3/4

OpenStudy (skullpatrol):

OK, you can solve it a few different ways, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I figured it out thanks :)

OpenStudy (skullpatrol):

Thanks for trying to understand :)